OpenStudy (nincompoop):

The specific heat of aluminum is 0.22 cal/g-ºC. What is the value of the specific heat in (a) kcal/kg-ºC, (b) J/kg-ºC? (Show your calculations)

OpenStudy (anonymous):

You can do it as a conversion problem using \[1 \ \text{cal}=1000 \ \text{kcal}\] \[4.2 \ J=1 \ \text{cal} \ \quad ;\text{or} \quad 4200 J=1 \text{kcal}\]

OpenStudy (anonymous):

Well that's wrong, sorry

OpenStudy (anonymous):

\[1000 \text{cal} = 1 \text{kcal}\]

OpenStudy (anonymous):

ie \[\left( \frac{.22 \ \text{cal}}{g \ ºC}\right) \left( \frac{1 \ \text{kcal}}{1000 \ \text{cal}} \right) \left( \frac{1000 g}{1 kg} \right) =\frac{.22 \ \text{kcal}}{kg ºC} \]

OpenStudy (nincompoop):

i know how ... laughing out loud I just can't copy and paste my math type into latex

OpenStudy (anonymous):

:P yah, equation editors aren't really compatible with each other...

OpenStudy (anonymous):

screen shots! ^_^

OpenStudy (nincompoop):

it used to allow me to do it but for some reason it is giving me errors this time

OpenStudy (anonymous):

it might be the degree sign for some reason? - that's what the masculine ordinal is

OpenStudy (nincompoop):

\[\begin{gathered} a) \hfill \\ \frac{{0.22cal}}{{g - C}} \times \frac{{1.00kcal}}{{1.00 \times {{10}^3}cal}} \times \frac{{1.00 \times {{10}^3}g}}{{1.00kg}} \cong \frac{{0.22kcal}}{{kg - C}} \hfill \\ \frac{\begin{gathered} b) \hfill \\ 0.22cal \hfill \\ \end{gathered} }{{g - C}} \times \frac{{4.186J}}{{1.00cal}} \times \frac{{1.00 \times {{10}^3}g}}{{1.00kg}} \cong \frac{{920J}}{{kg - C}} \hfill \\ \end{gathered} \] OMGGGGGGG

OpenStudy (nincompoop):

\[\begin{array}{l} a)\\ \frac{{0.22cal}}{{g - C}} \times \frac{{1.00kcal}}{{1.00 \times {{10}^3}cal}} \times \frac{{1.00 \times {{10}^3}g}}{{1.00kg}} \cong \frac{{0.22kcal}}{{kg - C}}\\ \frac{\begin{array}{l} b)\\ 0.22cal \end{array}}{{g - C}} \times \frac{{4.186J}}{{1.00cal}} \times \frac{{1.00 \times {{10}^3}g}}{{1.00kg}} \cong \frac{{920J}}{{kg - C}} \end{array}\]

OpenStudy (anonymous):

weird.... \hfill must be something that the latex package on the site doesn't use - it has things like \hspace and \hline. Dunno. Anyways, looks good methinks - on the first one though, I think you can use a definite equal sign, because the conversions are exact.

OpenStudy (nincompoop):

it was a different format or version of tex I used

OpenStudy (anonymous):

so nothing is working anymore, basically :P At least the site isn't crashing (for the moment)....

OpenStudy (nincompoop):

haha

OpenStudy (anonymous):

In any case, good luck with formatting fun!

OpenStudy (anonymous):

^_^